OpenStudy (anonymous):

Hi! Can anyone explain me how to calculate the activity of an oxidant/reductant? I have to calculate a silver/silver chloride electrode in physical chemistry.

OpenStudy (anonymous):

\[E=E _{0}+\frac{ RT }{ zF }\ln \frac{ a _{ox} }{ a _{red} }\] So I don't get how to derive a.

OpenStudy (vincent-lyon.fr):

In your case, the redox couple is \(Ag^+/Ag\) Half equation is \(Ag^++e^- \rightarrow Ag_{(s)}\) Which leads to \(E=E _{0}+\Large\frac{ RT }{ zF }\ln \Large \frac{ [Ag^+]}{ 1 }\)

OpenStudy (vincent-lyon.fr):

In the presence of chloride ions, \([Ag^+]=K_S/[Cl^-]\) Bythe way \(z=1\) here. After substituting [Ag+], you will get a workable equation.